1. Find the modulus and direction of the resultant force of the forces F1 and F2, the directions of which form the angle 90⁰ between them, and their modulus are

F1 = 12N and F2 = 8N.


1
Expert's answer
2021-02-07T19:10:52-0500

"F_t=\sqrt{12^2+8^2}=14.4\ (N)"


"\alpha=\tan^{-1}(\frac{12}{8})=56.3°" from the direction of force "F_2" . Answer